OpenStudy (anonymous):

f-1(3) when f(x) = 2 x minus 7, all over 3

OpenStudy (anonymous):

would i start this by putting 3 as X and making it 3 = 2 (3) - 7 / 3?

OpenStudy (anonymous):

solve \[\frac{2x-7}{3}=3\] for \(x\)

OpenStudy (anonymous):

so times 3 on both sides making it 2x - 7 = 9?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then add 7, divide by 2